# Global Energy Minimization for Simplex Mesh Optimization: A Radius Ratio Approach to Sliver Elimination

### v2

@@ 3.1 Fixed-point iteration algorithm
$\bm{G}_{F}=\begin{bmatrix}\bm{A}~&~\bm{B}_{2}~&~\bm{B}_{1}~\\ -\bm{B}_{2}~&~\bm{A}~&~\bm{B}_{0}~\\ -\bm{B}_{1}~&~-\bm{B}_{0}~&~\bm{A}\end{bmatrix}$
@@
$\begin{cases}\bm{A}\bm{X}+\bm{B}_{2}\bm{Y}+\bm{B}_{1}\bm{Z}=\bm{0}~\\ -\bm{B}_{2}\bm{X}+\bm{A}\bm{Y}+\bm{B}_{0}\bm{Z}=\bm{0}~\\ -\bm{B}_{1}\bm{X}-\bm{B}_{0}\bm{Y}+\bm{A}\bm{Z}=\bm{0}\end{cases}$
@@
$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}_{2}\bm{Y}^{old}-\bm{B}_{1}\bm{Z}^{old}-\bm{A}\bm{X}^{old}_{fix}~\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}_{2}\bm{X}^{old}-\bm{B}_{0}\bm{Z}^{old}-\bm{A}\bm{Y}^{old}_{fix}~\\ \bm{A}\bm{Z}_{free}^{new}=\bm{B}_{1}\bm{X}^{old}+\bm{B}_{0}\bm{Y}^{old}-\bm{A}\bm{Z}^{old}_{fix}\end{cases}$
@@
$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}\bm{Y}^{old}-\bm{A}\bm{X}^{old}_{fix}~\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}\bm{X}^{old}-\bm{A}\bm{Y}^{old}_{fix}\end{cases}$
@@ 3.2 PLBFGS Algorithm
$\delta=\begin{cases}0,\quad\text{if}~{}k\leq~m~\\ k-m,\quad\text{if}~{}k>m\end{cases};L=\begin{cases}k,\quad\text{if}~{}k\leq~m~\\ m,\quad\text{if}~{}k>m\end{cases};\bm{q}_{L}=\nabla F(\bm{x}_{k})$
@@ 3.3 PNLCG Algorithm
$\lambda_{k}$ is given by line search satisfying the strong Wolfe conditions with an initial guess $\lambda_{k}=1$. The scalar coefficient $\beta_{k}$ has multiple calculation formulas, see [16]. Notable $\beta_{k}$ formulas include: Polak-Ribière ($\beta_{k}^{PR}$[35], Hestenes–Stiefel ($\beta_{k}^{HS}$[18], Fletcher–Reeves ($\beta_{k}^{FR})$ [13], and Dai-Yuan($\beta_{k}^{DY}$[9]. We adopt the Polak-Ribière $\beta_{k}$ formula, defined as:
@@ 4 Experiment and Comparisons
We run all experiments on a laptop running Ubuntu 22.04. The CPU is an AMD Ryzen7 4800H(2.9GHz,8MB cache) and 16GB DDR4 RAM(3200MHz).
@@
Figure 5: For the L-shaped domain, the corner points and points on the domain boundary edges are fixed, while points on the boundary faces are allowed to move along the tangential directions of the faces. ODT optimization converges after 36 iterations.
@@
ODT 15 0.05079 106 19072 4.18
ODT 75 0.06376 58 18852 20.31
LBFGS 17(85) 0.10295 53 19784 7.95
PLBFGS 13(54) 0.18480 44 19784 5.84
NLCG 30(107) 0.16996 45 19784 9.89
PNLCG 16(69) 0.17120 43 19784 6.97
ODT+LBFGS 15+15(59) 0.28190 1 19072 4.18+3.64
ODT+PLBFGS 15+7(29) 0.30972 0 19072 4.18+2.16
ODT+PNLCG 15+7(38) 0.29942 1 19072 4.18+2.65
ODT+LBFGS 75+13(33) 0.31686 0 18852 20.31+2.04
ODT+PLBFGS 75+7(27) 0.32546 0 18852 20.31+2.01
ODT+PNLCG 75+8(34) 0.32572 0 18852 20.31+2.45
@@
ODT 15 0.01938 225 31846 13.48
ODT 36 0.06550 96 31371 31.49
LBFGS 14(109) 0.00556 6 32827 13.49
PLBFGS 24(145) 0.20427 10 32827 27.34
NLCG 10(82) 0.01216 12 32827 14.06
PNLCG 16(117) 0.12828 5 32827 22.22
ODT+LBFGS 15+22(91) 0.24053 4 31846 13.48+12.07
ODT+PLBFGS 15+12(61) 0.22503 4 31846 13.48+9.26
ODT+PNLCG 15+14(76) 0.19577 4 31846 13.48+11.50
ODT+LBFGS 36+12(51) 0.29676 1 31371 22.22+6.72
ODT+PLBFGS 36+8(32) 0.34373 0 31371 22.22+5.08
ODT+PNLCG 36+9(51) 0.31098 0 31371 22.22+7.57
@@
ODT 15 0.01119 843 68962 30.35
ODT 75 0.08138 188 66581 144.77
LBFGS 48(224) 0.17128 215 70105 83.62
PLBFGS 25(139) 0.19374 219 70105 56.42
NLCG 37(188) 0.09946 235 70105 36.26
PNLCG 16(88) 0.14231 217 70105 36.26
ODT+LBFGS 15+21(107) 0.22678 78 68962 30.35+30.85
ODT+PLBFGS 15+12(54) 0.21974 68 68962 30.35+17.92
ODT+PNLCG 15+11(61) 0.23623 66 68692 30.35+19.89
ODT+LBFGS 75+13(56) 0.24514 3 66581 144.77+15.74
ODT+PLBFGS 75+6(10) 0.28008 2 66581 144.77+4.39
ODT+PNLCG 75+7(35) 0.25638 3 66581 144.77+11.7
@@ 5 Conclusion and further work
In this paper, we propose a radius-ratio energy-based global mesh optimization algorithm that effectively suppresses sliver elements and improves tetrahedral mesh quality. We also design an efficient preconditioner to substantially enhance optimization efficiency. Our algorithm can operate independently or synergistically with existing mesh optimization methods. Benefiting from the preconditioner, it significantly improves mesh quality with minimal computational overhead when applied as a post-processing step to other optimization techniques.
Our algorithm is a vertex smoothing algorithm that does not include topological optimization. It performs well when mesh topology is already good. When our algorithm is combined with ODT algorithm, it can achieve good optimization results because ODT algorithm includes topology changes and effectively optimizes the mesh topology. When used alone, our algorithm can also optimize sliver elements, although the results are less effective than when combined with ODT. Furthermore, for surface boundaries, we currently use a fixed boundary point method, which also limits the optimization effect. Therefore, in future works, we plan to develop more effective boundary processing solution and add an edge/face exchange algorithm to further improve the algorithm, enhance the reliability of the algorithm, and extend this mesh optimization algorithm to anisotropic mesh. We also hope to optimize larger-scale mesh and further improve computational efficiency by using GPU parallel technology.

### v3

@@ 3.1 Fixed-point iteration algorithm
$\bm{G}_{F}=\begin{bmatrix}\bm{A}\nobreakspace&\nobreakspace\bm{B}_{2}\nobreakspace&\nobreakspace\bm{B}_{1}\nobreakspace\\ -\bm{B}_{2}\nobreakspace&\nobreakspace\bm{A}\nobreakspace&\nobreakspace\bm{B}_{0}\nobreakspace\\ -\bm{B}_{1}\nobreakspace&\nobreakspace-\bm{B}_{0}\nobreakspace&\nobreakspace\bm{A}\end{bmatrix}$
@@
$\begin{cases}\bm{A}\bm{X}+\bm{B}_{2}\bm{Y}+\bm{B}_{1}\bm{Z}=\bm{0}\nobreakspace\\ -\bm{B}_{2}\bm{X}+\bm{A}\bm{Y}+\bm{B}_{0}\bm{Z}=\bm{0}\nobreakspace\\ -\bm{B}_{1}\bm{X}-\bm{B}_{0}\bm{Y}+\bm{A}\bm{Z}=\bm{0}\end{cases}$
@@
$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}_{2}\bm{Y}^{old}-\bm{B}_{1}\bm{Z}^{old}-\bm{A}\bm{X}^{old}_{fix}\nobreakspace\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}_{2}\bm{X}^{old}-\bm{B}_{0}\bm{Z}^{old}-\bm{A}\bm{Y}^{old}_{fix}\nobreakspace\\ \bm{A}\bm{Z}_{free}^{new}=\bm{B}_{1}\bm{X}^{old}+\bm{B}_{0}\bm{Y}^{old}-\bm{A}\bm{Z}^{old}_{fix}\end{cases}$
@@
$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}\bm{Y}^{old}-\bm{A}\bm{X}^{old}_{fix}\nobreakspace\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}\bm{X}^{old}-\bm{A}\bm{Y}^{old}_{fix}\end{cases}$
@@ 3.2 PLBFGS Algorithm
$\delta=\begin{cases}0,\quad\text{if}~k\leq\nobreakspace m\nobreakspace\\ k-m,\quad\text{if}~k>m\end{cases};L=\begin{cases}k,\quad\text{if}~k\leq\nobreakspace m\nobreakspace\\ m,\quad\text{if}~k>m\end{cases};\bm{q}_{L}=\nabla F(\bm{x}_{k})$
@@ 3.3 PNLCG Algorithm
[…]
@@ 4 Experiment and Comparisons
We run all experiments on a laptop running Ubuntu 22.04. The CPU is an AMD Ryzen7 7735H(3.2GHz,16MB cache) and 16GB DDR5 RAM(5600MHz).
@@
Figure 5: This domain is an L-shaped domain with an empty ball domain inside. For the L-shaped domain, the corner points and points on the domain boundary edges are fixed, while points on the boundary faces are allowed to move along the tangential directions of the faces. ODT optimization converges after 36 iterations.
@@
ODT 15 0.05079 106 19072 2.70
ODT 75 0.06376 58 18852 13.05
LBFGS 29(110) 0.177782 42 19784 7.60
PLBFGS 14(60) 0.192391 43 19784 4.82
NLCG 28(125) 0.166748 41 19784 8.84
PNLCG 13(57) 0.181469 42 19784 4.36
ODT+LBFGS 15+15(59) 0.29190 1 19072 2.70+3.64
ODT+PLBFGS 15+7(29) 0.30972 0 19072 2.70+2.06
ODT+NLCG 15+15(62) 0.303464 0 19072 2.70+3.72
ODT+PNLCG 15+7(34) 0.30249 1 19072 2.70+2.33
ODT+LBFGS 75+13(33) 0.31686 0 18852 13.05+1.67
ODT+PLBFGS 75+7(25) 0.32531 0 18852 13.05+1.52
ODT+NLCG 75+14(56) 0.29968 0 18852 13.05+2.80
ODT+PNLCG 75+8(37) 0.312975 0 18852 13.05+2.12
@@
ODT 15 0.01938 225 31846 9.39
ODT 36 0.06550 96 31371 21.65
LBFGS 34(158) 0.22334 5 32827 20.78
PLBFGS 25(112) 0.24498 6 32827 17.08
NLCG 33(136) 0.209923 8 32827 17.69
PNLCG 18(81) 0.23201 4 32827 12.34
ODT+LBFGS 15+24(108) 0.21932 5 31846 9.39+10.67
ODT+PLBFGS 15+11(48) 0.23547 4 31846 9.39+5.86
ODT+NLCG 15+24(104) 0.20628 5 31846 9.39+10.38
ODT+PNLCG 15+16(83) 0.21572 4 31846 9.39+10.16
ODT+LBFGS 36+12(51) 0.29676 1 31371 21.65+5.03
ODT+PLBFGS 36+8(32) 0.34373 0 31371 21.65+3.94
ODT+NLCG 36+15(71) 0.34499 0 31371 21.65+6.95
ODT+PNLCG 36+8(37) 0.35575 0 31371 22.22+4.58
@@
ODT 15 0.01119 843 68962 20.14
ODT 75 0.08138 188 66581 97.46
LBFGS 42(192) 0.17838 220 70105 51.05
PLBFGS 28(153) 0.17877 216 70105 45.28
NLCG 33(148) 0.142602 219 70105 39.67
PNLCG 15(68) 0.16193 216 70105 20.53
ODT+LBFGS 15+25(112) 0.21412 70 68962 20.14+24.33
ODT+PLBFGS 15+22(103) 0.23569 68 68962 20.14+24.98
ODT+NLCG 15+24(113) 0.20821 69 68692 20.14+23.65
ODT+PNLCG 15+11(53) 0.20433 66 68692 20.14+12.89
ODT+LBFGS 75+13(56) 0.24514 3 66581 97.46+15.74
ODT+PLBFGS 75+6(10) 0.28008 2 66581 97.46+3.37
ODT+NLCG 75+14(59) 0.21657 5 66581 97.46+11.91
ODT+PNLCG 75+7(28) 0.28628 3 66581 97.46+6.74
@@ 5 Conclusion and further work
In this paper, we propose a radius-ratio energy-based global mesh optimization algorithm that effectively suppresses sliver elements and improves tetrahedral mesh quality. We also design an efficient preconditioner to substantially enhance optimization efficiency. Our algorithm can operate independently or synergistically with existing mesh optimization methods, such as ODT algorithm. Benefiting from the preconditioner, it significantly improves mesh quality with minimal computational overhead when applied as a post-processing step to other optimization techniques. In future works, we plan to develop more effective boundary processing solution and to further improve the algorithm, and extend this mesh optimization algorithm to anisotropic mesh. We also hope to optimize larger-scale mesh and further improve computational efficiency by using GPU parallel technology.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.